# Floquet breathers in a modulated nonlinear lattice

### v2

@@ III Effect of space-time modulation on the phonon dispersion relation
In this section, we present the main consequences of space-time modulation of the linear system (7). They are also valid for small displacements $u_{n}(t)$ in the nonlinearized model (9). The relatively detailed mathematical derivations based on Bloch theorem is included in Appendix B.
For $\delta=0$, that is, without modulation, the phonon dispersion relation (PDR) is given by $\omega_{q}^{2}=\omega_{0}^{2}+2\kappa(1-\cos(q))$ as a function of the wave number $q$ of the phonon. In this work, the minimum frequency $\omega_{0}=1$ and the maximum one $\sqrt{\omega_{0}^{2}+4\kappa}=1.01$ for $\kappa=0.052$. Numerically, the phonon spectrum can be obtained using the two-dimensional fast Fourier transform of $u_{n}(t)$ in space and time (XTFFT) as explained in App. B.1.
@@
Figure 3: Contour plot of the XTFFT after thermalization of the linear system with only time modulation. See App. B.1 for details. Parameters: $\kappa=0.1$, $\delta=0.075$, $h=0$, N=120 (Top panel) $\Omega=0.1$; (Bottom panel) $\Omega=0.01$
In App. B.3 it is deduced that to first order approximation the effect of time modulation alone is given by
@@
If there are both time and space modulation it is also deduced in App. B.4 that the phonon bands are given by:
@@
The derivations presented in Appendix B cannot predict the intensity of the various PDR. In App. B.2, it is also deduced the effect of symmetry breaking of the space invariance by only space modulation.
Figure 4: Contour plot of the XTFFT after thermalization of the linear system with both time and space modulation, together with the theoretical dispersion curves. See App. B.1 and B.4 for details. Parameters: $\delta=0.05$, $\Omega=0.05$, $\kappa=0.1$, $h=0.5$, N=120. Note that there is an area to the right where the harmonics are very close and become mixed.
@@ IV Breathers with space-time modulation
In this section we study the conditions for breather existence in a modulated system, first reviewing the theory of exact breathers and adapting it to space-time modulation and to just time modulation.
@@ IV.A Exact traveling breathers
The theory of moving breathers was developed by Aubry and coauthors in Refs.  Chen et al. (1996); Aubry and Cretegny (1998) and Flach Flach and Kladko (1999) among others. We will refer here to a formulation very similar to the latter, developed by some of the authors to interpret exact moving breathers in the $q-\omega$ space Archilla et al. (2019). Other articles related to that formulation are Ref. James and Sire (2005) with a rigorous mathematical treatment, and Gómez-Gardeñes et al. (2004) applied to the DNLS, where the concept of resonant lines (see below) appear. Here, we adapt the formulation in Ref. Archilla et al. (2019) to modulated systems.
Let us suppose that we have a space-time modulated system with phase $\phi=hn-\Omega t$ and velocity $V_{m}=h/\Omega$. An exact breather will be a solution of the form Flach and Kladko (1999):
@@
where the function $f$ is a localized function of its first argument and a $2\pi$ periodic function of the second, and $V_{b}$ is the velocity of the breather. If the breather is exact, there exists a minimal time $T_{F}$, called the fundamental time, and an integer $s$, called the step, such after a time $T_{F}$, the breather reproduces itself displaced a distance $s$ in lattice units. Then, $V_{b}=s/T_{F}$, and the fundamental frequency is defined as $\omega_{F}=2\pi/T_{F}$. The frequency $\omega_{i}$, is the moving frame frequency Archilla et al. (2019) also called internal frequency Flach and Kladko (1999).
@@
The resonant harmonics form straight lines within the $(q,\omega)$ space, called resonant lines. Their slope is $V_{b}$ and their intercept at the axis $q=0$ is $m\omega$, i.e, the moving frame frequency. Therefore, all the harmonics in a resonant line have the same frequency $m\omega_{F}$, with the integer $m$ indexing the lines.
The breather frequencies are inside one of those lines called the breather line with $m=m_{b}$. The intercept of the breather line with the axis $q=0$ is the moving frame frequency of the breather $\omega_{\text{b}}=m_{b}\omega_{F}$. All the harmonic waves in the same line have the same frequency in the moving frame and propagates with the same velocity $V_{b}$, which explains the persistence of the breather.
@@ IV.B Application to space-time harmonically modulated systems
For a space-time harmonically modulated system with a term $\cos(hn-\Omega t)$, it is also necessary that the value of the modulating phase $\Phi=hn-\Omega t$ is also identical modulo $2\pi$ after the translation $s$ and time change $T_{F}$, because if not, the forces will be different and the evolution of $u$ would be different.
@@
$h(n+s)-\Omega(t+T_{F})=hn-\Omega t+2\pi m_{m}\,\rightarrow$
$\Omega=\frac{hs+2\pi r}{T_{F}}\,\rightarrow\Omega=hV_{b}+m_{m}\omega_{F}\,.$ (14)
This means that the modulating wave is a resonant harmonic with index $m_{m}$ and moving frame frequency $m_{m}\omega_{F}$ and therefore the breather and modulating waves have commensurate moving frame frequencies.
@@
Lemma 2 For a harmonically, time-modulated system, a necessary condition for breather existence is that the breather frequency $\omega_{b}$ and the modulating frequency $\Omega$ are commensurate: $\omega_{\text{b}}/\Omega=m_{b}/m_{m}$.
Figure 5: (Top) For $\Omega=2\omega_{\text{b}}$, profile of the stable breather. (Bottom) Coordinates and energy evolution in a breather period $T_{b}$. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\omega_{\text{b}}=0.98$.
Figure 6: For $\Omega=2\omega_{\text{b}}$:(Top-left) Floquet multipliers for the non-autonomous system and (top-right) for the extended autonomous system. Blue circles have positive Krein signature, red diamonds negative one, and black squares correspond to zero Krein signature. The multipliers are identical except for the double +1 of the extended system. (Bottom panel) Eigenvector corresponding to the isolated eigenvalue, related with localized perturbations of the breather. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\omega_{\text{b}}=0.98$.
@@ IV.C Breather obtention in the time-modulated system
The most direct way is working in the real space with the Newton and shooting method. Trying some different localized initial conditions and observing an approximate periodical behavior with period close to $T_{b}$, the intended one, the initial conditions are used as a seed for the Newton method, where initial conditions corresponding to the exact $T_{b}$-periodic behavior. Exact means with a precision of $10^{-15}$, that is, very close, to the computer machine precision. The Newton method, a numerical application of the implicit function theorem, can be used as the system has no time invariance due to the time-periodic term in the potential, and a solution is unique in its neighborhood. The precision can be refined more by working in the space of the frequencies, if we consider time-reversible solutions, they can be written as $u_{n}(t)=\sum_{k=-k_{m}}^{k_{m}}z_{k,n}\exp(k\omega_{\text{b}}t)$. Due to $u_{n}(t)$ being real and time-symmetric, $z_{k,n}^{*}=z_{-k,n}$ and $z_{k,n}=z_{-k,n}$, and real. The $u_{n}(t)=z_{0,n}+\sum_{k=1}^{k_{m}}2z_{k,n}\cos(k\omega_{\text{b}}t)$, being 2zk,n2z{k,n} the inverse cosine fast Fourier cosine transform (idfc). Then, the Newton method can be coded in the space of the $k_{m}+1$ FFT components, finding the fftc components $z_{n,k}$. The value of $m=15$ provides excellent results.
@@ IV.D Stability of breathers in the time-modulated system
Let as suppose that the breather solution for a breather with $T_{b}=n_{T}T_{m}$, with $n_{T}$ an integer larger than 1, the $T_{m}$-periodic system is also $T_{b}$ periodic, and to analyze the stability we can construct the Floquet matrix integrating a small perturbation of the initial coordinates and momenta and observing the perturbation after $T=T_{b}$. For each changed coordinate we obtain a column of the Floquet matrix $F$. Its eigenvalues lie on the unit circle if the system is stable. For autonomous symplectic systems Aubry (2006) there are always two eigenvalues of $F$ at +1+1, corresponding to $u_{t}=\dot{u}$, the phase mode, as it represents that if $u(t)$ is a $T_{b}$ periodic solution, $u(t+dt)$ is also a solution, and the growth mode, indicating that around a solution there is another one with slightly different amplitude.
@@
$\ddot{u}_{n}-L_{n}u_{n}+\big(1+\delta_{2}\cos(\Omega t)\big)\frac{\displaystyle\partial V(u_{n})}{\displaystyle\partial u_{n}}=0\,$, (15)
@@
If $u_{b}(t)$ is a $T_{b}$-periodic solution of (15), and $u(t)=u_{b}(t)+\xi(t)$ a perturbed solution with $\xi(t)$ small, the evolution of $\xi$ is given by the Newton operator with zero eigenvalue, that is:
@@
However, it can be recovered within an autonomous extended system that includes the non-autonomous one, as shown below. For the different factors $n_{T}$, in $\Omega=n_{T}\omega_{\text{b}}$, we obtain different results that we will present later.

### v3

@@ III Effect of space-time modulation on the phonon dispersion relation
In this section, we present the main consequences of space-time modulation of the linear system (7). They are also valid for small displacements $u_{n}(t)$ in the nonlinear model (9). Appendix A explains the concept of thermalizaton and numerical experiments to observe the phonon dispersion relation (PDR). The relatively detailed mathematical derivations based on Bloch theorem are included in the rest of App. C.
For $\delta=0$, that is, without modulation, the phonon dispersion relation is given by $\omega_{q}^{2}=\omega_{0}^{2}+2\kappa(1-\cos(q))$ as a function of the wave number $q$ of the phonon. In most of this work, we use the physical value of the coupling parameter $\kappa=0.0052$, with the phonon frequencies between $\omega_{0}=1$ and $\sqrt{\omega_{0}^{2}+4\kappa}=1.010$. The exception is this section, where for some simulations we choose a larger value $\kappa=0.1$, with the phonon band between the frequencies 1 and 1.183, in order to have a good visualization of the shape of the original phonon band and the new bands that appear with the modulation. The parameters are specified in the captions.
Numerically, the phonon spectrum can be obtained using the two-dimensional discrete Fourier transform of $u_{n}(t)$ in space and time (XTDFT) as explained in App. C.1.
@@
Figure 3: Contour plot of the XTDFT after thermalization of the linear system with only time modulation. See App. C.1 for details. Parameters: $\kappa=0.1$, $\delta=0.075$, $h=0$, N=120 (Top panel) $\Omega=0.1$; (Bottom panel) $\Omega=0.01$.
In App. C.3, it is deduced that to first order approximation, the effect of time modulation alone is given by
@@
If there are both time and space modulation, it is also deduced in App. C.4 that the phonon bands are given by:
@@
The derivations presented in Appendix C cannot predict the intensity of the various PDR. In App. C.2, it is also deduced the effect of symmetry breaking of the space invariance by only space modulation.
Figure 4: Contour plot of the XTDFT after thermalization of the linear system with both time and space modulation, together with the theoretical dispersion curves. See App. C.1 and C.4 for details. Parameters: $\delta=0.05$, $\Omega=0.05$, $\kappa=0.1$, $h=0.5$, N=120. Note that there is an area to the right where the harmonics are very close and become mixed.
@@ IV Breathers with space-time modulation
In this section, we study the conditions for breather existence in a modulated system, first reviewing the theory of exact breathers and adapting it to space-time modulation and to only time modulation.
@@ IV.A Exact traveling breathers
The theory of moving breathers was developed by Aubry and coauthorsChen et al. (1996); Aubry and Cretegny (1998) and by Flach and Kladko Flach and Kladko (1999) among others. We will refer here to a formulation very similar to the latter, developed by some of the authors to interpret exact moving breathers in the $q-\omega$ space.Archilla et al. (2019) Other articles related to that formulation are Ref. James and Sire, 2005 with a rigorous mathematical treatment, and Ref. Gómez-Gardeñes et al., 2004 applied to the DNLS, where the concept of resonant lines (see below) appear. Here, we adapt the formulation in Ref. Archilla et al., 2019 to modulated systems.
Let us suppose that we have a space-time modulated system with phase $\phi=hn-\Omega t$ and velocity $V_{m}=h/\Omega$. An exact breather will be a solution of the form:Flach and Kladko (1999)
@@
where the function $f$ is a localized function of its first argument and a $2\pi$ periodic function of the second, and $V_{b}$ is the velocity of the breather. If the breather is exact, there exists a minimal time $T_{F}$, called the fundamental time, and an integer $s$, called the step, such after a time $T_{F}$, the breather reproduces itself displaced a distance $s$ in lattice units. Then, $V_{b}=s/T_{F}$, and the fundamental frequency is defined as $\omega_{F}=2\pi/T_{F}$. The frequency $\omega_{i}$, is the moving frame frequency Archilla et al. (2019) also called internal frequency.Flach and Kladko (1999)
@@
The resonant harmonics form straight lines within the $(q,\omega)$ space, called resonant lines. Their slope is $V_{b}$ and their intercept at the axis $q=0$ is $m\omega$, i.e, the moving frame frequency. Therefore, all the harmonics in a resonant line have the same moving frame frequency $m\omega_{F}$, with the integer $m$ indexing the lines.
The breather frequencies are inside one of those lines called the breather line with $m=m_{\text{b}}$. The intercept of the breather line with the axis $q=0$ is the moving frame frequency of the breather $\omega_{\text{b}}=m_{\text{b}}\omega_{F}$. All the harmonic waves in the same line have the same frequency in the moving frame and propagate with the same velocity $V_{b}$, which explains the persistence of the breather.
@@ IV.B Application to space-time harmonically modulated systems
For a space-time harmonically modulated system with a term $\cos(hn-\Omega t)$. It is also necessary that the value of the modulating phase $\Phi=hn-\Omega t$ is also identical modulo $2\pi$ after the translation $s$ and time change $T_{F}$, because, if not, the forces will be different and the evolution of $u$ would be different.
@@
$h(n+s)-\Omega(t+T_{F})=hn-\Omega t+2\pi m_{\text{m}}\,\Rightarrow$
$\Omega=\frac{hs+2\pi r}{T_{F}}\,\rightarrow\Omega=hV_{b}+m_{\text{m}}\omega_{F}\,$, (21)
with $m_{\text{m}}$ an integer. This means that the modulating wave is a resonant harmonic with index $m_{\text{m}}$ and moving frame frequency $m_{\text{m}}\omega_{F}$ and therefore the breather and modulating waves have commensurate moving frame frequencies.
@@
Note, that an exact moving breather is not periodic in time and has not a definite frequency, this condition applies to the moving frame frequencies and it is therefore different from Eq. 15, which reappears below.
@@
Lemma 2 For a harmonically, time-modulated system, a necessary condition for breather existence is that the breather frequency $\omega_{b}$ and the modulating frequency $\Omega$ are commensurate: $\omega_{\text{b}}/\Omega=m_{\text{b}}/m_{\text{m}}$.
Figure 5: (Top) For $\Omega=2\omega_{\text{b}}$, profile of the stable breather. (Bottom) Coordinates and energy evolution in a breather period $T_{\text{b}}$. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\omega_{\text{b}}=0.98$.
@@ IV.C Breather obtention in the time-modulated system
The most direct way is working in the real space with the Newton and shooting method. Testing some different localized initial conditions and observing an approximate periodical behavior with period close to $T_{\text{b}}$, the intended one. The initial conditions are used as a seed for the Newton method, in order to obtain the initial conditions corresponding to the exact $T_{\text{b}}$-periodic behavior. Exact means with a precision of $10^{-15}$, that is, very close to the computer machine precision. The Newton method, a numerical application of the implicit function theorem, can be used because the system has no time invariance due to the time-periodic term in the potential, and a solution is unique in its neighborhood. The precision can be refined further by working in the space of the frequencies. If we consider time-reversible solutions, they can be written as the inverse discrete Fourier transform $u_{n}(t)=\sum_{k=-k_{\text{m}}}^{k_{\text{m}}}z_{k,n}\exp(\textrm{i}k\omega_{\text{b}}t)$. Due to $u_{n}(t)$ being real and time-symmetric, $z_{k,n}^{*}=z_{-k,n}$ and $z_{k,n}=z_{-k,n}$ and real. Then, $u_{n}(t)=z_{0,n}+\sum_{k=1}^{k_{\text{m}}}2z_{k,n}\cos(k\omega_{\text{b}}t)$, and the Newton method can be coded to obtain the $k_{\text{m}}+1$ real coefficients $z_{k,n}$ for each site $n$. The value of $k_{\text{m}}=15$ provides excellent results.
@@ IV.D Stability of breathers in the time-modulated system
Figure 6: For $\Omega=2\omega_{\text{b}}$:(Top-left) Floquet multipliers for the non-autonomous system and (top-right) for the extended autonomous system. Blue circles have positive Krein signature, red diamonds negative one, and black squares correspond to zero Krein signature. The multipliers are identical except for the double +1 of the extended system. (Bottom panel) Eigenvector corresponding to the isolated eigenvalues, related with the internal modes of the breather. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\omega_{\text{b}}=0.98$.
Let as suppose a breather solution with $T_{\text{b}}=n_{T}T_{m}$, with $n_{T}$ an integer larger than 1, the $T_{m}$-periodic system is also $T_{\text{b}}$ periodic, and to analyze the linear stability we can construct the Floquet matrix integrating a small perturbation of the initial coordinates and momenta and observing the perturbation after $T=T_{\text{b}}$. For each changed coordinate or momentum we obtain a column of the Floquet matrix $F$. Its eigenvalues lie on the unit circle if the system is stable. For autonomous symplectic systems Aubry (2006) there are always two eigenvalues of $F$ at +1+1, corresponding to $u_{t}=\dot{u}$, the phase mode, as it represents the fact that if $u(t)$ is a $T_{\text{b}}$ periodic solution, $u(t+dt)$ is also a periodic solution, and the growth mode, indicating that around a solution there is another one with slightly different amplitude.
@@
$\ddot{u}_{n}-L_{n}u_{n}+\big(1+\delta_{2}\cos(\Omega t)\big)\frac{\displaystyle\partial V(u_{n})}{\displaystyle\partial u_{n}}=0\,$, ($2$)
@@
If $u_{b}(t)$ is a $T_{\text{b}}$-periodic solution of (22), and $u(t)=u_{b}(t)+\xi(t)$ a perturbed solution with $\xi(t)$ small, the evolution of $\xi$ is given by the Newton operator with zero eigenvalue, that is:
@@
It can also be checked numerically that the phase mode $\dot{u}_{\text{b}}(0),\ddot{u}_{\text{b}}(0)]^{t}=[0,\ddot{u}_{\text{b}}(0)]^{t}$ if not an eigenvector of the Floquet matrix $F$ with other eigenvalue. The isolated eigenvalues outside the phonon band correspond to eigenvectors with the same localization of the breather that are internal modes of the breather. In particular, when they have the same symmetry of the breather, they are called breathing modes.Johansson and Aubry (2000) Simulations with the eigenvectors added as a perturbation to the breather show that they produce a beating of the amplitude of the breather which justify their name.
@@
The double Floquet multiplier at +1 can be recovered within an autonomous extended system that includes the non-autonomous one, as shown below. For the different factors $n_{T}$, in $\Omega=n_{T}\omega_{\text{b}}$, we obtain different results that we will present later.
IV.E Extended autonomous symplectic system
The Hamiltonian structure of the dynamical equations, i.e., $\dot{u}_{n}=\partial H/\partial p_{n}$, $\dot{p}_{n}=-\partial H/\partial u_{n}$ is kept for the conjugate pairs of variables $u_{n}$, $p_{n}$, although the system is no longer autonomous. By defining $s=t$ or to be precise as (23) is $T_{\text{b}}$-periodic, as $s=\mathrm{rem}(t,T_{\text{b}})$, the remainder of $t/T_{\text{b}}$, i.e., $\mathrm{rem}(nT_{\text{b}}+t,T_{\text{b}})=t$ and $\mathrm{rem}(-nT_{\text{b}}-t,T_{\text{b}})=-t$, both for $n>0$ and $t>0$, so as $s$ is $T_{\text{b}}$ periodic. Then $\dot{s}=1$ as usual, and the system with the extra variable has become autonomous, but not symplectic as $s$ has no conjugate variable yet. The extra variable $h$ to make the system symplectic is obtained by defining a new Hamiltonian $K=H+h$, with $h$ the conjugate momentum of $s$ as $\dot{s}=\partial K/\partial h=1$ as previously obtained. Then, $\dot{h}=-\partial K/\partial s=-\partial H/\partial s=-\partial H/\partial t=-\textrm{d}H/\textrm{d}t$, or $h=E_{0}-H$, with $E_{0}$ the initial energy at $t=0$.Marthinsen and Owren (2016) The extended system is now symplectic and $\dot{u}_{b},\dot{s}]$ is now solution of (24) and being $T_{\text{b}}$ periodic, the Floquet multiplier +1+1 corresponding to the phase mode reappears. There is an extra +1+1 multiplier, because, symplectiness, implies that the multipliers are in pairs $\lambda,1/\lambda$, the extra mode corresponding to a solution with slightly different energy or growth mode.
The extended system reveals a hidden symplectic structure of the autonomous system with the implication that linear stability corresponds to all the Floquet multipliers being in the unit circle, but without the double multiplier of +1, which is recovered with the extended system. Figure 6-top-right shows these multipliers for $\Omega=2\omega_{\text{b}}$.
Note that these conclusions are valid both the cases $\Omega>wb$ and $\Omega\leq\omega_{\text{b}}$ studied in the next two sections taking due care of the time of integration. They are also valid for space-time modulated systems with trivial changes that we will detail elsewhere.
V Breathers with time modulation and $\Omega=n_{T}\omega_{\text{b}}$
We consider the nonlinear dynamical system (9) initially for the cases $\omega_{\text{b}}=\Omega/n_{T}$ ($T_{\text{b}}=n_{T}T_{\text{m}}$), and $\omega_{\text{b}}=n_{T}\Omega$ ($T_{\text{b}}=T_{\text{m}}/n_{T}$), $n_{T}$ an integer. The first one corresponds to the well known phenomenon of period doubling, tripling, etc., while the latter correspond to a breather that oscillates several times faster than the modulation. According to the deduction in Sec. IV, fractional frequencies are also possible and we will consider them in Sec. VII for different breather frequencies.
There is no dissipation of energy, but as the system is parametrically forced the energy is conserved only at discrete times separated by the larger of the two periods, the breather period $T_{\text{b}}$ and the modulating one $T_{m}$. This is shown in Fig. 5 for $T_{\text{b}}=2T_{m}$.
The dynamical equations (9) are invariant under the transformation $t\rightarrow t+T_{m}$, therefore, if a phonon $(q,\omega)$, i.e., $u=\exp(\textrm{i}[qn-\omega t])$ is a solution of the linearized equation, then also $\exp(\textrm{i}m\Omega t)u$ is also a solution, or, in other words, there is another phonon band with frequencies $\omega^{\prime}=\omega+m\Omega$, with $m$ a positive or negative integer. But there are also two phonon bands, with positive and negative frequencies. Then, if there is a phonon $(q,\omega)$, also there exists the phonon $(q,\Omega-\omega)$. As the unmodulated phonon band for $\kappa=0.052$ is in $\omega_{0}=1,\omega_{\text{max}}=1.01]$, for $\Omega\simeq 2$, a new phonon band appears in $\Omega-\omega_{\text{max}}\simeq 0.98,\Omega-\omega_{0}\simeq 1]$. The phonon bands in the modulated system change position, and $\Omega=2\omega_{\text{b}}=1.96$, but those bands are very close with the possibility of interfering, depending on the exact value of $\omega_{\text{b}}$ and actual modulated phonons. This problem does not exist for $\Omega\geq 3$ as the band $\Omega-\omega$, for $\omega$ in the unmodulated phonon band is much more separated.
For $\Omega=2\omega_{\text{b}}$, and changing slightly $\omega_{\text{b}}$, there is an interval where the two phonon bands coincide and then the system becomes stimulated by $\Omega$ and the variables diverge, as it is also shown in Sec. VII. A dissipation term can be added, but as this is a very special case, we prefer not to modify the system and simply avoid those combination of frequencies.
V.A Period doubling $\Omega=2\omega_{\text{b}}$
The subject of time-modulation is of recent interest in photonic topological insulators. For example, Floquet solitons, for which the intensity repeats after each driving period, were found experimentally in optical waveguide arrays with Kerr nonlinearity Mukherjee and Rechtsman (2020) and period-doubled Floquet solitons were observed in a photonic topological insulator formed by an array of waveguides with a honeycomb pattern.Mukherjee and Rechtsman (2023) For these solutions, the intensity period is doubled. The same happens in a helically modulated honeycomb lattice as well and a sinusoidally driven SSH lattice. These references show that the period-doubling phenomenon can appear in many different physical systems of interest and, in particular, photonic systems are ideal experimental devices to observe the connection between time modulation, topology and nonlinearity.
Period doubling also appears in our system. In Fig. 5-bottom we can see the evolution of the breather center and its neighboring sites, also the evolution of the energy, which is $T_{m}$-periodic and therefore also $T_{\text{b}}$-periodic. This is a consequence of $u_{n}(T_{\text{b}}/2)=-u_{n}(0)$, $p_{n}(T_{\text{b}}/2)=-p_{n}(0)=0$, the energy being a function of $u_{n}^{2}$ and $\cos(\Omega t)=1$ at values of $t$ that are integer multiples of $T_{b}/2$. There is no net gain or loss of energy every $T_{\text{b}}/2$ with a relative precision of $10^{-9}$ after 100 breather periods using a symplectic integrator.Sanz-Serna and Calvo (1994)
The Floquet eigenvalues are also pictured in Fig. 6-top-left. The eigenvalues corresponding to the phonons form a pair of complex conjugate bands width frequencies $\omega_{n}=\omega_{\text{b}}+\theta_{n}/T_{\text{b}}$, as they have the same Krein signature,Aubry (2006) they form a structurally stable subset. There is also a close but separate eigenvalue with a localized eigenvector shown in Fig. 6-bottom, that corresponds to the breathing mode, a type of internal breather mode as explained in Sec. IV.D . It has the opposite Krein signature of the phonons, meaning it can collide with the phonon with the closest frequency leaving the unit circle, which indicates that the breather is unstable and another solution exists. As shown in Sect. VII, in some cases the new solution corresponds to a breather in quadrature, i.e., with a difference of phase of $\pi/2$, with the modulation function.